Reports 1-1 of 1 Clear search Modify search
AdV-TCS (Ring Heater)
berni, gherardini, nardecchia, - 22:43 Sunday 25 February 2024 (63398) Print this report
SR RH tuning attempt

This morning, we continued the SR RH tuning started yesterday. With the ITF LN3 for 2 hrs, Francesco changed the SR RH power from 7 W to 9 W.

09.07 UTC: SR RH increase up to 9 W (V= 26.23 V)

Immediately after, the ITF unlocked. The ITF relocked at LN2 at 10.57 UTC.

I asked Francesco to leave the ITF in LN2 to monitor the DCP during the thermal transient.

Between 13 UTC and 14.30 UTC, the ITF experienced 3 equal unlocks just after the OMC lock.

In the meanwhile, at each lock, the 112 MHz sideband appeared lower, thus the last step of SR RH power has been reverted decreasing the RH power to 7 W (see fig.1)

14.38 UTC:  SR RH voltage reverted at 23.13 V (P=7 W)

After that, the lock acquisition has been stopped in CARM null 1F while Fabio manually aligned the SR mirror to survive to the transient.

At 17.22 UTC, the lock acquisition has been resumed but the ITF unlocked during ‘ACQUIRE LN1’.

The subsequent lock acquisition was successful and ITF relocked at LN2 at 17.57 UTC. Then, the ITF unlocked at 18.38 UTC.

In all collected LN2, the lock ended after ~1hr with a fast increase of B1p. 

After this lock,  we were not able to relock anymore.

We collected several unlocks during ‘ACQUIRE LN1’ that seem all equal (Fabio's entry). 

The positive/negative effect of this SR RH working point (at 7 W) needs to be checked.

Images attached to this report
Search Help